OpenStudy (anonymous):

find the slope and y - intercept 5x + 4y = 12

OpenStudy (saifoo.khan):

your turn.

OpenStudy (anonymous):

-5/4m and 3=b

OpenStudy (anonymous):

um is this right?? \[5x + 4y = 12\] \[5(0) + 4y = 12\] \[4y/4 = 12/4\] \[y = 3\]

OpenStudy (anonymous):

and zero slope?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

that is the y intercept. you did that part correct.

OpenStudy (anonymous):

-5/4m should say -5/4=m m being the slope

OpenStudy (saifoo.khan):

correct.